OpenStudy (anonymous):

To reduce 16 Oz of a 18% solution to a 10% solution how much distilled water should be added

OpenStudy (littlebird):

18 % = mass solute x 100/ 16 mass solute =2.88 ounces 10 = 2.88 x 100/X X = mass solution = 28.8 ounces mass water that must be added = 28.8 - 16=12.8 ounces

OpenStudy (anonymous):

Thank you I always like to double check how I solve things